# Classical and quantum beam dynamics simulation of the RF photoinjector test bench

### v3

@@ Abstract
We present beam-dynamics simulations for an S-band RF photoinjector test bench under development at the Joint Institute for Nuclear Research, aimed at producing high-quality electron beams and enabling future generation of relativistic vortex electrons with quantized orbital angular momentum (OAM). Simulations of the 1.5-cell photogun under the presently available RF gradient of 45 MV/m demonstrate stable bunch formation at low charge ($Q=0.63$ pC), where space-charge forces are weak and the transverse emittance is dominated by RF-induced correlations. Optimization of the injection phase and cathode-region solenoid yields a robust emittance compensated regime with a final normalized emittance of $3.2\ \pi\cdot\mathrm{mm\cdot mrad}$. To assess prospects for accelerating vortex electron beams, we additionally model the quantum evolution of single-electron Laguerre–Gaussian wave packets. The results show that multi-MeV acceleration suppresses free-space spreading of the electron packet and preserves the packet’s initial OAM structure, indicating that the test bench provides suitable conditions for forthcoming experimental studies of relativistic vortex electrons.
@@ 1 Introduction
The generation of relativistic vortex electron beams – quantum states with a helical phase front and a quantized orbital angular momentum projection (OAM) – is attracting growing interest across accelerator, atomic, and nuclear physics, electron microscopy, quantum optics, and foundation of quantum mechanics [Knyazev18, Bliokh17, Karlovets21, Ivanov22]. To date, vortex electrons have been generated only in the sub-MeV range, up to kinetic energies of $\sim 300$ keV of transmission electron microscopes [Uchida10, Verbeeck10, Blackburn14]. Extending such structured states into the multi-MeV regime relevant for photoinjectors and linacs remains an outstanding challenge, yet promises significant impact, including new tools for nuclear and hadronic structure studies and potential alternatives to spin-polarized beams [Ivanov22, SchattschneiderPRB12, Wu19, An25].
@@
It is important to distinguish the genuine vortex beams of single particles whose wave functions are characterized by a helical phase structure $\exp(i\ell\varphi)$ and a quantized OAM projection $\ell\hbar$ [Bliokh17] from beams with modified transverse emittance ratios – such as flat and round beams demonstrated at Fermilab [Danilov96, Brinkmann02, Sun04, Piot06]. These latter beams can posses a so-called extrinsic angular momentum, which is a property of the beam as a whole, so a quantum state of each particle in the beam is not defined, whereas the former quantum beams of single electrons posses the so-called intrinsic orbital angular momentum, the property of each electron, complementary to spin [Bliokh17, Karlovets21, Ivanov22, Floettmann20].
@@
One feasible route towards high-energy vortex electrons is the transfer of OAM from twisted ultraviolet light to electrons during photoemission [Karlovets21, Pavlov24, Kazinski25]. Whereas the generation of pure vortex states with a definite angular momentum projection and its vanishing quantum uncertainty is challenging due the finite temperature, space-charge effects, and other factors [Karlovets21, Floettmann20, Kazinski25], the generation of a superposition of vortex states with an uncertainty less than the mean OAM value of the beam seems feasible when the OAM is very high, in our case up to $\ell=64\hbar$, and the bunch charge is low so the space-charge effects can be mitigated.
@@
Realizing this mechanism requires a photoinjector capable of producing low-emittance beams with well-controlled transverse profiles and rapid acceleration, ensuring that the initial OAM-imprinted phase structure of the emitted electron is preserved during the early acceleration. To this end, a dedicated S-band RF photoinjector test bench is being developed at the Laboratory of Nuclear Problems, Joint Institute for Nuclear Research (JINR), combining a 1.5-cell RF gun [Nikiforov18] and a high-power UV laser system [Gacheva14]. Along with relativistic vortex electrons, there are plans at the Institute of Modern Physics in China to put a twist on relativistic ions [An25].
@@
RF photoinjectors are widely used as high-brightness electron sources at facilities such as LCLS [Alley99], the European XFEL [Ferrario01], and SuperKEKB [Cahill14], owing to their ability to quickly suppress space-charge effects via high-gradient acceleration [Giribono23, Maltseva24]. At JINR, there exists a LINAC facility, which is a linear accelerator of electrons to the energy of 200 MeV and plans to come to 400 MeV in 2027 [Nozdrin25], it operates with a thermionic DC gun and a multi-stage RF bunching system [Nozdrin21], producing $\sim 1$22 ps bunches. Replacing it with an RF photoinjector enables direct generation of 10 ps bunches in single-bunch operation, with the significantly improved transverse beam quality and precise control over the charge and transverse profile – critical prerequisites for producing structured quantum states.
@@
In this work, we present beam-dynamics simulations of the bunch generation and acceleration in the JINR photoinjector using realistic field maps. We demonstrate the implementation of a Carlsten-type emittance-compensation scheme. Although the emittance does not decrease by an order of magnitude, the solenoidal lens stabilizes the transverse emittance at a low value and suppresses RF-induced phase-space distortions, establishing a robust operating point for subsequent vortex-beam experiments.
@@ 2 Test bench design
The 1.5-cell S-band RF gun operates in the $\pi$-mode at 2856 MHz [Samofalova21]. Its internal geometry and photocathode placement (see Fig. 3) are optimized for rapid extraction and emittance-compensating beam formation. The UV laser enters through the coaxial channel, which simultaneously serves as the RF coupling element and the electron-beam exit port. The on-axis accelerating field (see Fig. 3) was obtained from CST Microwave Studio eigenmode simulations [CST].
@@
The emittance-compensation solenoid prototype (see Fig. 5) was manufactured at JINR. It has an inner radius of 100 mm, outer radius of 140 mm, length of 100 mm, and contains 216 copper turns in six layers. The axial field profile, measured with 5-mm resolution at 20 A, is shown in Fig. 5. These measured values are used directly in the beam-dynamics simulations.
@@
3 Beam dynamics simulation
3.1 Definition of initial parameters
Beam dynamics simulations were carried out with $\mathrm{ASTRA}$ [ASTRA], which couples a 3D particle-in-cell space-charge solver (in cylindrical symmetry) to high-order Runge–Kutta tracking through RF-cavity and magnetic fields. The key inputs are the bunch charge (from photocathode quantum efficiency and laser energy), mean initial energy and its spread, initial normalized emittance and spatial/momentum distributions, and the RF phase at photoemission.
@@ 3.1 Definition of initial parameters
Using a copper photocathode with a quantum efficiency of $QE\sim 10^{-5}$ [Xiang15], this yields an emitted charge of approximately
@@
The photoemission model follows the Dowell–Schmerge approach [Dowell09], assuming Fermi–Dirac statistics and Gaussian spatial distributions. For an effective copper work function of $\phi_{\rm eff}=4.3\ \mathrm{eV}$ [Karkare15], the resulting mean kinetic energy and energy spread of emitted electrons lie in the sub-eV range, consistent with standard metallic photocathodes. With the UV laser rms beam size set to $\sigma_{x,y}=2\ \mathrm{mm}$ (equal to the initial RMS beam size), the corresponding intrinsic normalized emittance is
@@
Figure 7: Simulated dependence of the final normalized transverse emittance on the injection RF phase $\varphi_{0}$. The operational working region $\varphi_{0}\approx 180^{\circ}$.
For realistic simulations, apertures corresponding to the internal structure of the RF gun were added (see Fig. 3): two apertures with an inner diameter of d=21.8mmd=21.8\ $\mathrm${mm} and a length of 10 mm, simulating the resonator cell walls, as well as an aperture with an inner diameter of d=14mmd=14\ $\mathrm${mm} and a length of 190 mm, simulating the coaxial input for impedance matching and RF power injection into the resonator cells. Simulations without the external solenoidal field demonstrate that most of the particles from the generated beam are lost within the first 25 cm from the cathode inside the coaxial input.
At the present low bunch charge (0.63pC0.63\ $\mathrm${pC}), the emittance is thus governed mainly by RF-induced transverse correlations imposed by the cavity geometry and mode structure, rather than by space-charge forces. This field configuration explains the observed monotonic phase dependence of the emittance. The operational working point of the test bench, chosen to maximize the energy gain under the available RF gradient, lies at $\varphi_{0}\approx 180^{\circ}$. In this region the field on the cathode is not maximal, which inevitably increases the extraction-induced emittance, but it provides the highest attainable beam energy at the current stage of commissioning. The results of this scan define the initial parameters used in the subsequent beam-dynamics simulations (Table 1).
@@
Normalized emittance $1.07\,\pi\,\mathrm{mm\,mrad}$
@@
Total beamline length 1.568 m
3.2 Beam dynamics without emittance compensation
This process is identical for both simulation scenarios, with and without the emittance-compensating solenoid, as it depends solely on the amplitude of the RF field and the proper injection phase, and is independent of external magnetic fields.
@@ 3.2 Beam dynamics without emittance compensation
Fig.9 shows the simulated mean kinetic energy of the bunch along the RF photogun beamline. A two-peak structure arises from the axial field in the S-band RF photogun, with rapid acceleration to $\overline{W_{\rm kin}}=1.9$ MeV at the gun exit, followed by a plateau. High-gradient injection quickly suppresses space-charge driven emittance growth once $\gamma\gg 1$, although early nonlinear distortions remain significant. Longitudinally, the bunch length $\sigma_{z}$ (see Fig.9) increases rapidly during acceleration due to the velocity–energy correlation, as faster electrons at the head outpace slower ones at the tail; beyond the cavity, the expansion continues more gradually under residual space-charge forces.
The simulated energy evolution is defined by the presently available RF power. Consequently, the accelerating gradient is reduced to 45MV/m45\ $\mathrm${MV/m}, resulting in a final kinetic energy of 1.9MeV1.9\ $\mathrm${MeV} in Fig. 9. The reduced gradient also increases the relative importance of space-charge forces during the first few millimeters after the emission.
Because the space-charge effects are weak in this regime, they can be neglected in the subsequent analysis. This is particularly important for the quantum-mechanical considerations in Sec. 4, since the small Coulomb forces do not significantly perturb the phase structure or the orbital-angular-momentum content of the evolving electron wave packet.
3.3 Emittance compensation using a solenoidal lens
A solenoidal lens is installed near the cathode (see Fig. 5) to counteract early nonlinear space-charge effects. The on-axis field profile used in the simulations (see Fig. 5) was optimized via iterative scans, yielding a peak of $B_{z}=125$ mT at $z=85$ mm downstream of the photocathode.
Fig. 13 shows the resulting RMS transverse beam size. Immediately after the RF cavities, the solenoid collimates the envelope, stabilizing $\sigma_{x,y}$ along the beamline. The normalized emittance (see Fig. 13) initially rises due to the non-zero solenoidal field at the cathode, which imparts extrinsic angular momentum onto the beam. In addition, the combination of the solenoidal field with the transverse RF fields of the 1.5-cell $\pi$-mode gun modifies the evolution of the transverse phase space and suppresses the natural cancellation of RF-induced correlations that occurs in the zero-solenoid case [Floettmann15].
Figure 12: Simulated RMS transverse beam size along the beamline with the compensation solenoid. The size stabilizes soon after the RF cavities, indicating effective collimation.
Figure 13: Simulated normalized transverse emittance along the beamline with solenoid compensation. Stabilizes near $3.2\ \pi\cdot\mathrm{mm}\cdot\mathrm{mrad}$ downstream.
@@ 3.3 Emittance compensation using a solenoidal lens
In addition, Fig. 15 shows the two-dimensional transverse phase space of the beam. The $xx^{\prime}$ projection exhibits an almost ideal phase-space ellipse, indicating stable transverse dynamics and good beam quality, even at relatively large transverse emittance values. Since the simulation corresponds to on-axis beam transport through the solenoid, an identical phase-space distribution is expected in the $yy^{\prime}$ plane.
@@ 4.1 Vortex electrons
In general, a vortex electron beam is a free-electron state whose wave function can be represented in terms of an LG mode, an exact non-stationary solution of the Schrödinger equation and a close analogue of optical vortex modes [Allen92]. Its transverse wave function can be written as [Karlovets21]
@@
One mechanism for generating vortex electrons in an RF photoinjector at an accelerator facility is laser-driven emission using structured light. Theory shows that an atom placed on the axis of an optical vortex transfers a well-defined OAM projection to the emitted electron, and this behaviour persists for ensembles [Pavlov24, Kazinski25]. Alternative schemes based on electron diffraction [Uchida10, Verbeeck10, Blackburn14] are impractical in such setups.
@@ 4.2 Quantum evolution of a single-electron wave packet in the accelerating field
To analyze the generation of relativistic vortex electrons carrying OAM, we follow the quantum evolution of a single-electron wave packet during acceleration. As a convenient observable we consider the mean-squared transverse radius $\langle\rho^{2}\rangle(\langle z\rangle)$, which for an LG wave packet is related to the transverse rms width via [Karlovets21]
@@
The general law of transverse quantum spreading for free (field-less) wave packets [Karlovets21] can be written either as a function of the mean propagation distance $\langle z\rangle$ or of the corresponding time of flight $t(\langle z\rangle)$:
@@
Here $z_{R}$ is the Rayleigh length and $t_{d}$ is the diffraction time [Karlovets21],
@@
In the far-field regime, $\langle z\rangle\gg z_{R}$, Eq. (4) reduces to a simple relation, often referred to as a van Cittert–Zernike–type theorem [Karlovets21], between the measured packet width at the detector and the initial width near the source,
@@
For an initial rms transverse radius $\sqrt{\langle\rho^{2}\rangle(0)}=1\ \mathrm{nm}$, typical for an electron packet near the emission region [Karlovets21, Ehberger15], with $n=3$ and $\ell=1$ (OAM $\ell\hbar$) and a drift length of 30cm30\ $\mathrm${cm}, Eq. (4) predicts free-space spreading to an rms transverse width of about 0.8m0.8\ $\mathrm${m} for an initial kinetic energy of 0.3eV0.3\ $\mathrm${eV}. Next, we include acceleration.
@@
where for acceleration of electrons along +z+z one has $E<0$ and hence $qE>0$, so that $t(z)>0$. The transverse second moment at a fixed plane $z$ is then obtained by evaluating Eq. (4) at this $t(z)$. For typical DC-gun this results in a $\sim 550$-fold reduction of the final transverse width compared to free-space propagation (see Fig. 17), in good agreement with recent relativistic calculations [z6j7-grs2].
@@
In the S-band photoinjector electrons are accelerated in a standing-wave RF field. Along the axis, where the packet propagates, the longitudinal electric field can be written as [Wiedemann15, Chao20, Lee21]

### v4

@@ Abstract
We present beam-dynamics simulations for an S-band RF photoinjector test bench under development at the Joint Institute for Nuclear Research, aimed at producing high-quality electron beams and enabling future generation of relativistic vortex electrons with a quantized orbital angular momentum (OAM). Simulations of the 1.5-cell photogun are performed assuming an RF gradient of 45 MV/m, which, in accordance with our simulations with CST Studio, corresponds to the currently achieved input RF power of 3 MW. At low charge ($Q=0.63$ pC), stable bunch formation is obtained, with weak space-charge effects and transverse emittance dominated by RF-induced correlations. Optimization of the injection phase and cathode solenoid results in a robust emittance-compensated regime with a final normalized emittance of $2.08\ \pi\cdot\mathrm{mm\cdot mrad}$. To assess prospects for accelerating vortex electron beams, we additionally model the quantum evolution of single-electron Laguerre–Gaussian wave packets. The results show that multi-MeV acceleration suppresses free-space spreading of the electron packet and preserves the packet’s initial OAM structure, indicating that the test bench provides suitable conditions for forthcoming experimental studies of relativistic vortex electrons.
@@ 1 Introduction
The generation of relativistic vortex electron beams – quantum states with a helical phase front and a quantized orbital angular momentum projection (OAM) – is attracting growing interest across accelerator, atomic, and nuclear physics, electron microscopy, quantum optics, and foundation of quantum mechanics [27, 5, 24, 22]. To date, vortex electrons have been generated only in the sub-MeV range, up to kinetic energies of $\sim 300$ keV of transmission electron microscopes [41, 42, 4]. Extending such structured states into the multi-MeV regime relevant for photoinjectors and linacs remains an outstanding challenge, yet promises significant impact, including new tools for nuclear and hadronic structure studies and potential alternatives to spin-polarized beams [39, 44, 21, 20, 22, 3].
@@
It is important to distinguish the genuine vortex beams of single particles whose wave functions are characterized by a helical phase structure $\exp(i\ell\varphi)$ and a quantized OAM projection $\ell\hbar$ [5] from beams with modified transverse emittance ratios – such as flat and round beams demonstrated at Fermilab [10, 6, 40, 36]. These latter beams can posses a so-called extrinsic (non-quantized) angular momentum, which is a property of the beam as a whole, so a quantum state of each particle in the beam is not defined, whereas the former quantum beams of single electrons posses the so-called intrinsic (quantized) orbital angular momentum, the property of each electron, complementary to spin [5, 15, 24, 22].
@@
One feasible route towards high-energy vortex electrons is the transfer of OAM from twisted ultraviolet light to electrons during photoemission [24, 35, 25]. Whereas the generation of pure vortex states with a definite angular momentum projection and its vanishing quantum uncertainty is challenging due the finite temperature, space-charge effects, and other factors [15, 24, 25], the generation of a superposition of vortex states with an uncertainty less than the mean OAM value of the beam seems feasible when the OAM is very high, in our case up to $\ell=64\hbar$, and the bunch charge is low so the space-charge effects can be mitigated.
@@
Realizing this mechanism requires a photoinjector capable of producing low-emittance beams with well-controlled transverse profiles and rapid acceleration, ensuring that the initial OAM-imprinted phase structure of the emitted electron is preserved during the early acceleration. To this end, a dedicated S-band RF photoinjector test bench is being developed at the Laboratory of Nuclear Problems, Joint Institute for Nuclear Research (JINR), combining a 1.5-cell RF gun [32] and a high-power UV laser system [18]. Along with relativistic vortex electrons, there are plans at the Institute of Modern Physics in China to put a twist on relativistic ions [3].
@@
RF photoinjectors are widely used as high-brightness electron sources at facilities such as LCLS [2], the European XFEL [14], and SuperKEKB [7], owing to their ability to quickly suppress space-charge effects via high-gradient acceleration [19, 30]. At JINR, there exists a LINAC facility, which is a linear accelerator of electrons to the energy of 200 MeV and plans to come to 400 MeV in 2027 [34], it operates with a thermionic DC gun and a multi-stage RF bunching system [33], producing $\sim 1$22 ps bunches. Replacing it with an RF photoinjector enables direct generation of 10 ps bunches in single-bunch operation, with the significantly improved transverse beam quality and precise control over the charge and transverse profile – critical prerequisites for producing structured quantum states.
@@
In this work, we present beam-dynamics simulations of bunch generation and acceleration in the JINR photoinjector using realistic electromagnetic field maps. We demonstrate that the solenoidal lens provides effective transverse-envelope control in the low-charge regime. Although the normalized emittance does not decrease below its intrinsic value, the solenoid stabilizes the transverse phase-space evolution and suppresses RF-induced distortions, establishing a robust and reproducible operating point for subsequent vortex-beam experiments.
@@ 2 Test bench design
The 1.5-cell S-band RF gun operates in the $\pi$-mode at 2856 MHz [38]. Its internal geometry (see Fig. 3) is optimized for rapid extraction and emittance-compensating beam formation. The UV laser enters through the coaxial channel, which simultaneously serves as the RF coupling element and the electron-beam exit port. The on-axis accelerating field (see Fig. 3) was obtained from CST Microwave Studio eigenmode simulations [9].
@@
The emittance-compensation solenoid prototype (see Fig. 5) was manufactured at JINR. It has an inner radius of 100 mm, outer radius of 140 mm, a mechanical length of 100 mm, and contains 216 copper turns arranged in six layers.The axial magnetic-field profile was measured at a current of 20 A with a longitudinal step of 5 mm. The measured distribution is shown in Fig. 5. Because the discrete measurement resolution and experimental constraints make it difficult to resolve a perfectly smooth field decay toward the edges, the measured data were numerically interpolated and slightly smoothed to obtain a continuous field profile suitable for beam-dynamics simulations. The resulting profile preserves the measured peak field and overall integral strength, while providing a physically consistent description of the fringe fields extending beyond the mechanical length of the coil. This interpolated axial field map is used in the simulations.
@@
3 Classical beam dynamics and operating regime
Beam dynamics simulations were carried out with $\mathrm{ASTRA}$ [16], which combines a 3D particle-in-cell space-charge solver (in cylindrical symmetry) with high-order Runge–Kutta tracking in external RF and magnetic fields. The principal input parameters are the bunch charge (determined by photocathode quantum efficiency and laser pulse energy), the mean initial energy and its spread, the normalized transverse emittance, the initial spatial and momentum distributions, and the RF phase at photoemission.
@@ 3 Classical beam dynamics and operating regime
Using a copper photocathode with a quantum efficiency of $QE\sim 10^{-5}$ [45], this yields an emitted charge of approximately
@@
The photoemission model follows the Dowell–Schmerge approach [11], assuming Fermi–Dirac statistics and Gaussian spatial distributions. For an effective copper work function of $\phi_{\rm eff}=4.3\ \mathrm{eV}$ [23], the resulting mean kinetic energy and energy spread of emitted electrons lie in the sub-eV range, consistent with standard metallic photocathodes. With the UV laser rms beam size set to $\sigma_{x,y}=2\ \mathrm{mm}$ (equal to the initial RMS beam size), the corresponding intrinsic normalized emittance is
@@
To account for realistic transport, apertures corresponding to the internal RF-gun geometry were included (Fig. 3). Without an external solenoidal field, most particles are lost within the first 25 cm from the cathode inside the coaxial input structure.
At the present low bunch charge (0.63pC0.63\ $\mathrm${pC}), collective space-charge effects are significant only within the first few millimeters after emission and rapidly diminish as the beam becomes relativistic. Consequently, the transverse emittance evolution is governed primarily by RF-induced correlations and by magnetization effects associated with the solenoidal field, rather than by Coulomb forces.
The operational working point of the test bench, selected to maximize the achievable energy gain at the available RF gradient, corresponds to $\varphi_{0}\approx 180^{\circ}$ (see Appendix A). Although the accelerating field at the cathode is not maximal at this phase, this setting yields the highest exit energy under commissioning conditions. The parameters used in subsequent simulations are summarized in Table 1.
@@
Normalized emittance 1.07  $\pi\cdot$ $m$ $\cdot$ mrad
@@
Total beamline length 0.3 m
Fig. 7 presents the simulated mean kinetic energy along the beamline. The characteristic two-peak structure reflects the axial field distribution of the 1.5-cell S-band RF gun. The beam exits the gun with $\overline{W_{\rm kin}}=1.9$ MeV after which the energy remains constant in the drift. The accelerating gradient is presently limited to 45MV/m45\ $\mathrm${MV/m} by the available RF power.
The longitudinal rms bunch length $\sigma_{z}$ (Fig. 7) increases rapidly during acceleration due to the velocity–energy correlation established in the RF field. As the head of the bunch gains slightly higher velocity than the tail, ballistic expansion occurs. Downstream of the cavity, the growth continues gradually under residual energy spread, while space-charge effects are already strongly suppressed by relativistic acceleration.
@@
Because of the low bunch charge, Coulomb interactions do not significantly distort the downstream phase-space structure. This is particularly relevant for the quantum-mechanical considerations discussed in Sec. 4, where preservation of the phase structure is essential.
Figure 8: Simulated RMS transverse beam size $\sigma_{x,y}$ along the beamline in the presence of the compensation solenoid. After a short matching region, the envelope remains nearly constant at $\approx 2.3$ $m$, indicating stable axisymmetric transport.
Figure 9: Simulated normalized transverse emittance along the beamline with solenoidal field at the cathode. The emittance increases inside the solenoid due to beam magnetization and subsequently converges to a constant intrinsic value of $2.08\ \pi\cdot\mathrm{mm\cdot mrad}$ downstream.
A compensation solenoid is installed near the cathode (Fig. 5) to control the transverse envelope during the low-energy stage. The simulated on-axis field profile (Fig. 5) was optimized iteratively, yielding a peak field of $B_{z}=125$ mT at $z=85$ mm.
Figure 9 shows the rms transverse beam size along the beamline. After a short adjustment region immediately downstream of the gun, the envelope remains nearly constant, $\sigma_{x}=\sigma_{y}=2.3$ mm, demonstrating stable, magnetically controlled transport without noticeable radial expansion. The symmetry between $x$ and $y$ confirms the predominantly axisymmetric dynamics.
The normalized transverse emittance evolution is shown in Fig. 9. Inside the solenoid region, the emittance exhibits a pronounced increase, reaching approximately $\varepsilon_{n,x,y}\approx 5.6\ \pi\cdot\mathrm{mm\cdot mrad}$, which reflects the magnetization acquired by the beam in the presence of a non-zero axial magnetic field at the cathode. The solenoidal field introduces transverse correlations associated with angular momentum, temporarily increasing the projected emittance. The interplay between the solenoidal field and the transverse RF fields of the 1.5-cell $\pi$-mode gun modifies the natural cancellation of RF-induced correlations that occurs in the zero-solenoid case [17].
Downstream of the solenoid, the projected emittance decreases and converges to a stationary value of $\varepsilon_{n,x,y}=2.08\ \pi\cdot\mathrm{mm\cdot mrad}$. In the absence of strong nonlinear forces, this value remains constant throughout the remainder of the beamline, indicating that the intrinsic transverse phase-space area is preserved during linear transport.
@@
Figure 10 shows the transverse phase-space distribution at the RF-gun exit. The $xx^{\prime}$ projection forms an approximately elliptical distribution with no pronounced filamentation, consistent with controlled envelope evolution. Due to axisymmetry, an identical structure is expected in the $yy^{\prime}$ plane. The overall picture confirms that, at the present low charge, the beam dynamics is dominated by RF-induced correlations and controlled solenoidal magnetization, while space-charge effects remain subdominant after the initial acceleration stage.
@@ 4.1 Vortex electrons
In general, a vortex electron beam is a free-electron state whose wave function can be represented in terms of an LG mode, an exact non-stationary solution of the Schrödinger equation and a close analogue of optical vortex modes [1]. Its transverse wave function can be written as [24]
@@
One mechanism for generating vortex electrons in an RF photoinjector at an accelerator facility is laser-driven emission using structured light. Theory shows that an atom placed on the axis of an optical vortex transfers a well-defined OAM projection to the emitted electron, and this behaviour persists for ensembles [35, 25]. Alternative schemes based on electron diffraction [41, 42, 4] are impractical in such setups.
@@ 4.2 Quantum evolution of a single-electron wave packet in the accelerating field
To analyze the generation of relativistic vortex electrons carrying OAM, we follow the quantum evolution of a single-electron wave packet during acceleration. As a convenient observable we consider the mean-squared transverse radius $\langle\rho^{2}\rangle(\langle z\rangle)$, which for an LG wave packet is related to the transverse rms width via [24]
@@
The general law of transverse quantum spreading for free (field-less) wave packets [24] can be written either as a function of the mean propagation distance $\langle z\rangle$ or of the corresponding time of flight $t(\langle z\rangle)$:
@@
Here $z_{R}$ is the Rayleigh length and $t_{d}$ is the diffraction time [24],
@@
In the far-field regime, $\langle z\rangle\gg z_{R}$, Eq. (4) reduces to a simple relation, often referred to as a van Cittert–Zernike–type theorem [24], between the measured packet width at the detector and the initial width near the source,
@@
For an initial rms transverse radius $\sqrt{\langle\rho^{2}\rangle(0)}=1\ \mathrm{nm}$, typical for an electron packet near the emission region [24, 13], with $n=3$ and $\ell=1$ (OAM $\ell\hbar$) and a drift length of 30cm30\ $\mathrm${cm}, Eq. (4) predicts free-space spreading to an rms transverse width of about 0.8m0.8\ $\mathrm${m} for an initial kinetic energy of 0.3eV0.3\ $\mathrm${eV}. Next, we include acceleration.
@@
where for acceleration of electrons along +z+z one has $E<0$ and hence $qE>0$, so that $t(z)>0$. The transverse second moment at a fixed plane $z$ is then obtained by evaluating Eq. (4) at this $t(z)$. For typical DC-gun this results in a $\sim 550$-fold reduction of the final transverse width compared to free-space propagation (see Fig. 11), in good agreement with recent relativistic calculations [31].
@@
In the S-band photoinjector electrons are accelerated in a standing-wave RF field. Along the axis, where the packet propagates, the longitudinal electric field can be written as [43, 8, 28]